OpenStudy (anonymous):

Use finite approximations to estimate the area under the curve f(x) = sin(x) between 0 and x=pi using an upper sum with four rectangles of equal length

OpenStudy (anonymous):

urrr reimann sum D:

OpenStudy (anonymous):

Yes! but I am not sure how to do this with sin(x)!

OpenStudy (anonymous):

Is this right? dx = (b-a)/n dx = pi/4 area1 = (pi/4)(sin(pi/4) + area2 = (pi/4)(sin(2pi/4)+ area3 = (pi/4)(sin(3pi/4) + area4 = (pi/4)sin(pi) ????????????? and if it is, now what? I know sin(pi/2) = 1 and sin(pi) = 0 but I don't know the other 2...